# MSGNav: Unleashing the Power of Multi-modal 3D Scene Graph for Zero-Shot Embodied Navigation

### v2

@@ 4.2 Main Experimental Results
We will show the main comparison results with other state-of-the-art methods on the Goat-Bench and OVON datasets. For clarity, we report only the primary results here, and detailed results will be provided in the Appendix.
4.2.1 HM3D-OVON
@@
BCRL [bcrl] × 8.0 2.5
DAgger [dagger] × 10.2 4.7
DAgRL [dagger] × 18.3 7.9
RL [RL] × 18.6 7.5
VLFM [vlfm] 35.2 19.6
Uni-NaVid [uninavid] × 39.5 19.8
TANGO [tango] 35.5 19.5
DAgRL+OD [ovon] × 37.1 19.9
MTU3D [mtu3d] × 40.8 12.1
MSGNav (Ours) 48.3 27.0
Table 1: Experiments on the “Val Unseen” split of HM3D-OVON.
Table 1 demonstrates that our proposed MSGNav significantly outperforms all other methods on both SR and SPL metrics in open-vocabulary navigation. Notably, while achieving the highest SR of 48.3%, MSGNav also exhibits an exceptionally high SPL, surpassing the previous state-of-the-art training-based method MTU3D [mtu3d] by 7.5% in SR and 14.9% in SPL. This shows that our approach effectively balances accuracy and efficiency in the navigation exploration process for unseen scenarios.
4.2.2 Goat-Bench
The results in Table 2 also show the outstanding performance of our MSGNav in tackling multimodal goals, lifelong navigation task settings. MSGNav achieves the best performance with 52.0% SPL and 29.6% SPL compared to both previously training-based and train-free methods. It surpasses the previous state-of-the-art training-based method MTU3D [mtu3d] by 4.8% SR and 2.1% SPL, demonstrating its superior navigation capabilities. These results highlight the effectiveness of our multi-modal scene graph in tackling multi-modal lifelong navigation tasks.
Method Training-free SR $\uparrow$ SPL $\uparrow$
SenseAct-NN Monolitic [goatbench] × 12.3 6.8
Modular CLIP on Wheels [goatbench] 16.1 10.4
Modular GOAT [goatbench] 24.9 17.2
SenseAct-NN Skill Chain [goatbench] × 29.5 11.3
VLMnav [vlmnav] 20.1 9.6
DyNaVLM [dynavlm] 25.5 10.2
3D-Mem [3dmem] 28.8 15.8
TANGO [tango] 32.1 16.5
MTU3D [mtu3d] × 47.2 27.7
@@ 4.2.2 Goat-Bench
Table 3: Component ablation experiment across the first episode of each scene on the “Val Unseen” split of GOAT-Bench. The first row without any module, which represents our baseline model 3D-Mem [3dmem] results. “VVD”, “AVU”, and “CRV” represent the Visibility-based Viewpoint Decision module, Adaptive Vocabulary Update module, and Closed-loop Reasoning and Verification module.
@@
Table 4: Scene graph experiment across the first episode of each scene on the “Val Unseen” split of GOAT-Bench. “Node-only” indicates Concept-graph [Conceptgraphs] without object relation edges. “Traditional graph” indicates Concept-graph [Conceptgraphs].
@@ 4.3.2 Advantage of M3DSG
Table 4 shows the experimental results comparing MSGNav using Node‑only, Concept‑graph [Conceptgraphs], and our M3DSG. Concept‑graph outperforms Node‑only by 4.4% in SR and 1.5% in SPL, especially for Language and Image goals, underscoring the value of relationship edges. M3DSG further boosts these categories with a total of 3.8% SR and 4.3% SPL, showing its effectiveness in scene understanding. It is worth noting that node-only performs well on category goals, mainly because such tasks rely on concise object sets rather than complex contextual information.
@@ 5 Conclusion
Limitations and future work. Despite these advantages of MSGNav, scene graph-based methods still face low inference efficiency due to the latency of VFMs and VLMs, suggesting future work on faster graph construction and inference for real‑time deployment. Additionally, while the “last-mile” problem has been mitigated by the VVD module, it is not fully resolved. Exploring reinforcement learning approaches, e.g. active target recognition, to further solve this challenge is also a way worthy of attention.

### v3

@@ 4.2 Main Experimental Results
We will show the main comparison results with other state-of-the-art methods on the Goat-Bench [16] and HM3D-ObjNav [31] benchmark.
4.2.1 Goat-Bench Benchmark
The results in Table 1 also show the outstanding performance of our MSGNav in tackling multimodal goals, lifelong navigation task settings. MSGNav achieves the best performance with 52.0% SPL and 29.6% SPL compared to both previously training-based and train-free methods. It surpasses the previous state-of-the-art training-based method MTU3D [55] by 4.8% SR and 2.1% SPL, demonstrating its superior navigation capabilities. These results highlight the effectiveness of our multi-modal scene graph in tackling multi-modal lifelong navigation tasks.
@@ 4.2.1 Goat-Bench Benchmark
SenseAct-NN Monolitic [16] × 12.3 6.8
Modular CLIP on Wheels [16] 16.1 10.4
Modular GOAT [16] 24.9 17.2
SenseAct-NN Skill Chain [16] × 29.5 11.3
VLMnav [10] 20.1 9.6
DyNaVLM [15] 25.5 10.2
3D-Mem [41] 28.8 15.8
TANGO [56] 32.1 16.5
MTU3D [55] × 47.2 27.7
@@
4.2.2 HM3D-ObjNav Benchmark
Table 1 demonstrates the superior zero-shot navigation capabilities of our MSGNav on multi-modal open-vocabulary tasks in GOAT-Bench, we further evaluate it on the widely used HM3D-ObjNav benchmark.
As shown in Table 1, MSGNav achieves a state-of-the-art Success Rate (SR) of 63.0%, outperforming the previous best training-free method, Apexnav [51], by 3.4%. While our Success Path Length (SPL) of 31.4% is slightly lower than that of Apexnav, this may be caused by the VVD module prioritizing high-visibility viewpoints over absolute shortest paths, which slightly increases the path length but significantly boosts the overall success rate.
Method Training-free SR $\uparrow$ SPL $\uparrow$
ProcTHOR-ZS [8] × 20.2 8.2
ZSON [27] × 25.5 12.6
PSL [37] × 42.4 19.2
L3MVN [46] 48.7 23.3
GOAT [4] × 50.6 24.1
ESC [54] 39.2 22.3
VLFM [44] 52.4 30.3
OpenFMNav [20] 52.5 24.1
SG-Nav [42] 54.0 24.9
Uni-goal [43] 54.5 25.1
Trihelper [50] 56.5 25.3
Apexnav [51] 59.6 33.0
MSGNav (Ours) 63.0 31.4
Table 2: Experiments on the HM3D-ObjNav benchmark.
@@ 4.2.2 HM3D-ObjNav Benchmark
Table 3: Component ablation experiment across the first episode of each scene on the “Val Unseen” split of GOAT-Bench. The first row without any module, which represents our baseline model 3D-Mem [41] results. “VVD”, “AVU”, and “CRV” represent the Visibility-based Viewpoint Decision module, Adaptive Vocabulary Update module, and Closed-loop Reasoning and Verification module.
@@
Table 4: Scene graph experiment across the first episode of each scene on the “Val Unseen” split of GOAT-Bench. “Node-only” indicates Concept-graph [11] without object relation edges. “Traditional graph” indicates Concept-graph [11].
@@ 4.3.2 Advantage of M3DSG
Table 4 shows the experimental results comparing MSGNav using Node‑only, Concept‑graph [11], and our M3DSG. Concept‑graph outperforms Node‑only by 4.4% in SR and 1.5% in SPL, especially for Language and Image goals, underscoring the value of relationship edges. M3DSG further boosts these categories with a total of 3.8% SR and 4.3% SPL, showing its effectiveness in scene understanding. It is worth noting that node-only performs well on category goals, mainly because such tasks rely on concise object sets rather than complex contextual information.
@@ 5 Conclusion
Limitations and future work. Despite these advantages of MSGNav, scene graph-based methods still face low inference efficiency due to the latency of VFMs and VLMs, suggesting future work on faster graph construction and inference for real‑time deployment.
Acknowledgements. This work was supported in part by the National Natural Science Foundation of China (No.42571514).
Supplementary Material
6 Implement Details
This section outlines the implementation details of our approach, including computational resource, hyper-parameters, and the vision foundation models (VFMs) used for constructing M3DSG.
6.1 Computational Resource
Our approach primarily employs a Vision-Language Model (VLM) for inference. Consequently, when utilizing API calls to closed-source models, it places only modest demands on local GPU computational resources. All experiments were conducted using NVIDIA A800-SXM4-40GB graphics cards, with each card requiring approximately 8 GB to 10 GB of VRAM to perform VFMs inference. This configuration ensures stable performance while maintaining cost-effectiveness in resource usage.
Meanwhile, inference with open-source large models such as GPT-OSS  [^2] GPT-OSS: https://github.com/openai/gpt-oss and LLAVA  [^3] LLAVA: https://github.com/haotian-liu/LLaVA is also feasible. However, this configuration demands additional GPU memory that scales in proportion to the parameter size of open-source large models.
6.2 Hyper-parameters
The details of the hyperparameters of MSGNav are shown in Table 6. Additionally, we maintain consistency with 3D-Mem and Conceptgraph for other parameters in the 3D scene graph construction.
Hyperparameters Value
Radii set: $\mathbf{R}$ [0.5, 0.75] (m)
Camera height: $h$, 1.5 (m)
Success threshold: $d$ 0.25 (m)
Samples number: $K$ 20
Adjacency distance: $\theta$ 3.5 (m)
Obstruction distance: $\tau$ 0.05 (m)
Top‑k related objects: $|\textbf{O}^{rel}|$ 20
Hfov of camera 120
Image resolution 1280x1280
Image prompt resolution 512x512
Table 6: The details of hyperparameters of MSGNav.
6.3 VFMs for Constructing M3DSG
As outlined in the main paper, the 3D scene reconstruction pipeline primarily incorporates the following Visual Feature Modules (VFMs). First, YOLO-World [6] is applied for object detection and room recognition. Next, SAM [18] is utilized to generate semantic masks. Finally, CLIP [29] is employed to extract visual feature representations of the objects.
An example of visualized perception results is shown as Fig. 6, illustrating the outputs obtained after processing with YOLO-World and SAM for open-vocabulary object detection and semantic segmentation.
Figure 6: A visualization example of perception results during constructing the M3DSG.
And the detailed information of VFMs is shown in Table 7.
VFM Version Size
Yolo-World [6] YOLOv8x-world 141.11 MB
SAM [18] SAM-Large 1.16 MB
CLIP [29] ViT-H-14-quickgelu/dfn5b 12.20 GB
Table 7: The detailed information of VFMs for constructing M3DSG.
VLM Overall (2669) Object Category (991) Language (856) Image (822)
(MSGNav) Success Rate SPL Success Rate SPL Success Rate SPL Success Rate SPL
GPT-4o [14] 51.97 29.56 53.38 29.98 47.43 25.08 54.99 33.71
Qwen-VL-Max [1] 52.79 30.79 56.10 31.53 45.56 27.18 56.33 33.67
Table 8: Experiments of our MSGNav method using different VLMs on the “Val Unseen” split of GOAT-Bench. The number following each category represents the sample size.
7 More Experimental Results
7.1 Additional Analysis on HM3D-OVON
To further investigate the foundational perception and optimization stability of the proposed MSGNav, we conduct supplementary experiments on the HM3D-OVON dataset [45]. For this analysis, we specifically report the results on the “Val Seen” split to evaluate the model’s performance when navigating within familiar environment distributions, which serves as a benchmark for the model’s upper-bound efficiency and basic semantic mapping capabilities.
Method Training-free SR $\uparrow$ SPL $\uparrow$
BCRL [33] $\times$ 20.2 8.2
DAgger [33] $\times$ 18.1 9.4
DAgRL [33] $\times$ 41.3 21.2
RL [39] $\times$ 39.2 18.7
VLFM [44] 35.2 18.6
Uni-NaVid [49] $\times$ 41.3 21.2
DAgRL+OD [45] $\times$ 35.8 21.2
MTU3D [55] $\times$ 55.0 23.6
MSGNav (Ours) 48.3 27.0
Table 9: Comparison of navigation performance on the Val Seen split of HM3D-OVON. This experiment evaluates the model’s optimization effectiveness and navigation efficiency within the training distribution.
As presented in Table 9, we compare MSGNav with several state-of-the-art training-based and training-free methods on the seen environments. While the training-based method MTU3D [55] achieves a higher Success Rate (SR) of 55.0% by leveraging extensive environment-specific learning, MSGNav significantly outperforms all methods in terms of Success weighted by Path Length (SPL), reaching 27.0%.
This result is particularly noteworthy as MSGNav is a training-free approach. The superior SPL indicates that even in seen scenarios, our method generates more efficient and purposeful navigation paths compared to models that might overfit to specific trajectories. These supplementary results confirm that MSGNav maintains robust basic navigation logic and high path efficiency, which provides a solid foundation for the zero-shot generalization capabilities demonstrated in our primary experiments (Table 1).
7.2 Quantitative experiments on Goat-bench
We present the detailed experimental results in Table 2 of the main paper, showing the detailed performance of MSGNav across different categories with various VLM. As shown in the table 8, MSGNav demonstrates strong adaptability across different VLMs instead of relying on the specific model. MSGNav achieves comparable results on both state-of-the-art VLMs, Qwen-VL-Max and GPT-4o. It is worth noting that the results on Qwen-VL-Max are slightly better than those on GPT-4o. This is primarily due to the costly token fees of GPT-4o, our prompts were primarily implemented based on Qwen-VL-Max, and they may be better suited for Qwen-VL-Max. However, the excellent results achieved on GPT-4o without any prompt adjustments demonstrate that the system does not overly rely on prompt engineering.
8 Analysis of the Exploration
8.1 Additional Details of VVD Module
Figure 7: Qualitative Visualization of the Visibility-based Viewpoint Decision (VVD) Module. The central plot shows a top-down Bird’s-Eye View (BEV) of the agent’s spatial environment. Red stars indicate potential targets, while red circles mark the Ground Truth (GT) target viewpoints. The VVD module evaluates candidate viewpoints (numbered 1, 2, and 3), calculating a visibility score based on ray-casting to the target point cloud. Viewpoint 1 achieves the ’Best’ visibility score (0.97), and its corresponding first-person observation provides a clear, unoccluded view of the red refrigerator. Viewpoint 2 receives a ’Medium’ score (0.83), capturing a partial, side-angle view. Viewpoint 3 receives the ’Worst’ score (0.05), as its line of sight is heavily occluded by other environmental structures.
We provide an example illustrating how the VVD module selects optimal viewpoints. As shown in Fig. 7, candidate viewpoints located closer to the ground-truth (GT) viewpoints receive higher visibility scores. In this example, the highest-scoring viewpoint (Viewpoint 1, score 0.97) is positioned directly in front of the target, offering an unobstructed field of view that clearly captures the target object.
By contrast, the medium-scoring viewpoint (Viewpoint 2, score 0.83) is positioned on the side of the target, thereby capturing only partial visual information. Notably, Viewpoint 3 (score 0.05) is severely occluded by environmental structures, which may be less discernible purely from a simplified 2D bird’s-eye view (BEV) representation. In such cases, the VVD module correctly identifies the occlusion by computing 3D line-of-sight, consistently assigning low visibility scores to these obstructed viewpoints.
To further clarify the mechanics of the VVD module, we expand upon the operations detailed in Algorithm 2. The algorithm systematically evaluates line-of-sight occlusion in 3D space to identify the optimal navigation endpoint.
  1. 1.
    Candidate Generation (Lines 1-4): The algorithm first computes the centroid $\mathbf{c}_{\bar{o}}$ of the localized target’s point cloud $\mathcal{PC}_{\bar{o}}$. It then generates a set of $K$ candidate viewpoints in a concentric circle (or multiple concentric circles defined by radii set $\mathbf{R}$) around this centroid. Crucially, it filters this set to retain only viewpoints $\mathbf{V}_{c}$ that reside in traversable free space, ensuring the agent can physically reach them.
  2. 2.
    Ray Construction (Line 7): For a given candidate viewpoint $\mathbf{v}_{i}$ and a specific point $\mathbf{p}$ belonging to the target object, the algorithm mathematically defines a ray segment $\mathcal{Q}(\mathbf{v}_{i},\mathbf{p})$. This segment connects the viewpoint and the target point. To account for the physical volume of the agent and potential sensor noise, the parameter $\tau$ introduces a safety margin at both ends of the ray.
  3. 3.
    Occlusion Evaluation (Line 8): The core visibility condition $\mathcal{E}(\mathbf{v}_{i},\mathbf{p})$ evaluates whether the defined ray $\mathcal{Q}$ is obstructed by the broader scene point cloud $\mathcal{PC}$. A target point $\mathbf{p}$ is considered visible if and only if the minimum distance between any point $\mathbf{q}$ on the ray and any point $\mathbf{s}$ in the environment point cloud is strictly greater than or equal to the obstruction margin $\tau$.
  4. 4.
    Visibility Scoring (Lines 9-13): The algorithm computes an aggregate visibility score $S_{\mathbf{v}_{i}}$ for the candidate viewpoint. This score represents the ratio of target points $\mathbf{p}\in\mathcal{PC}_{\bar{o}}$ that satisfy the visibility condition $\mathcal{E}$. The viewpoint achieving the highest ratio is selected as the optimal navigation target $\mathbf{v}_{best}$, guaranteeing the agent finishes its trajectory with a robust, unoccluded view of the target.
8.2 Visualization about VVD Module
Figure 8: An example of viewpoint decision by the VVD module. This shows the BEV view of the point cloud in the scene and the corresponding viewpoint information.
Figure 9: This visualization illustrates the process of an agent exploring and navigating to a target described by natural language—“refrigerator in the kitchen. It is located next to the kitchen cabinet and the worktop”—in an indoor environment. It is composed of three parts from top to bottom: the Top-down Map sequence, the Observation sequence (first-person visual inputs), and the legend, which respectively display the global exploration state, the agent’s local visual perception, and the meaning of each visual element.
Figure 10: Visualization of the VLM reasoning process based on M3DSG. The agent successfully locates the target ”berber carpet” by leveraging M3DSG Nodes to match the object candidates (Object 13) and Multi-modal Edges (images) to verify complex spatial relationships, such as the carpet’s placement near the refrigerator (Object 11) and its position in the bottom right corner of the frame. This accurate grounding leads to an efficiently successful navigation task.
We provide a real-world example illustrating how the VVD module selects optimal viewpoints. As shown in Fig. 8, candidate viewpoints located closer to the ground-truth (GT) viewpoints, marked by the purple triangle, receive higher visibility scores. In this example, the highest-scoring viewpoint is almost coincident with the GT viewpoint and is positioned directly in front of the target (top), offering an unobstructed field of view.
By contrast, the previous viewpoint selection was farther from the GT viewpoint and positioned on the side of the target, thereby capturing only partial visual information. Notably, no GT viewpoint exists at the rear (lower side) of the target due to occlusion caused by point cloud data (green dots) at the same height within the scene, which may be less discernible in the bird’s-eye view (BEV) representation. In such cases, the VVD module consistently assigns low visibility scores (0.1–0.6) to these occluded viewpoints.
8.3 Visualization about exploration
As shown in Fig. 9, we present the natural language-guided target navigation by visualizing the agent’s process of exploring and localizing the target described as “refrigerator in the kitchen. It is located next to the kitchen cabinet and the worktop” to intuitively demonstrate its closed-loop reasoning.
The agent starts at the predefined Start Position and begins to explore adjacent areas. At this stage, the target refrigerator is still unseen (indicated by a red dot). The agent continuously moves along the Frontier (the boundary between explored and unexplored areas, marked by purple arrows) to uncover more regions. Its movement trajectory is recorded as the Exploration path (light gray line), and its Current Position (blue dot) is updated dynamically as it navigates through different rooms. As the agent ventures into the kitchen area, the Observation sequence captures the visual input of the refrigerator adjacent to the kitchen cabinet and worktop. Simultaneously, in the Top-down Map, the target marker changes from a red dot (Target (unseen)) to a green dot (Target (seen)), and the agent reaches the Final Position (green triangle), completing this hard navigation task.
This visualization comprehensively demonstrates how the agent uses the multi-modal scene graph to sequentially reason and localize the target object in an unknown environment. With the help of our multi-modal scene graph, agents can gradually explore and locate targets even when they are extremely distant. This further validates the effectiveness of our multi-modal scene graph for reasoning in zero-shot embodied navigation tasks.
Furthermore, to explicitly illustrate the critical role of our proposed M3DSG in the granular decision-making process, we visualize the Vision-Language Model (VLM) reasoning process in Fig. 10. In this episode, the agent is tasked with finding a target based on a complex spatial description: ”herber carpet is the room. It is located near the rack, freezer, and refrigerator towards the bottom right corner of the frame”. As the exploration proceeds ($t=1\dots T$), the M3DSG incrementally updates. When matching the target, the VLM relies on the M3DSG Nodes to identify the candidate, explicitly noting that the described object matches Object 13. More importantly, it heavily depends on the Multi-modal Edges (stored images) to verify fine-grained spatial relations. The reasoning process highlights that specific images (particularly Image 1, 2, and 5) show a large red carpet positioned near a refrigerator (Object 11), a shelf, and a freezer. Additionally, the visual edges confirm the carpet’s placement near the bottom right corner of the frame, perfectly aligning with the language description. By combining node attributes and visual edge evidence, the agent confidently concludes that Object 13 is the correct match. Ultimately, this accurate multimodal reasoning allows the agent to efficiently succeed in the task, reaching the target viewpoint at a remarkably close distance of 0.127m.
Figure 11: Example illustrating the use of a Vision-Language Model (VLM) for multimodal scene graph reasoning.
8.4 Analysis of multi-modal scene Graph reasoning
We present an example illustrating the use of a Vision-Language Model (VLM) for multimodal scene graph reasoning. As shown in Fig. 11, when given the query “plant that is located near the window glass and radiator”, the conventional textual description in the scene graph failed to include information pertinent to the target. In contrast, the VLM, utilizing the supplementary visual information embedded in the scene graph, successfully identified the target matching the query. The model then enriched the scene graph through vocabulary supplementation and updates. As a result, the task was completed with a Success weighted by Path Length (SPL) of 89.79, requiring only about 3 m of exploration length. This demonstrates that retaining images within the multimodal scene graph substantially improves exploration efficiency.
Figure 12: Prompt 1 for “top-k object nodes selection in KSS module”. The placeholders {…} will be replaced by the corresponding information. The gray-highlighted text represents the information of compressed M3DSG.
9 VLM Reasoning and Prompt
Our MGSNav method leverages Vision-Language Model (VLM) reasoning in three scenarios during the navigation process:
  • Top-k object nodes selection in KSS (As shown in Fig. 12) — Within the Key Scene Selection (KSS) module, the VLM is applied to infer the top-$k$ target-related nodes from the compressed scene graph.
  • Exploration inference — The KSS-compressed key scene subgraph is processed by the VLM to identify either the target node or the exploration frontier. This reasoning is performed in two sequential steps:
    1. 1.
      Find target. If the target is present in the 3D scene graph, input the scene graph information to identify the target ID. (As shown in Fig. 13)
    2. 2.
      Explore frontier. If absent, input the frontier image to determine the exploration frontier image ID. (As shown in Fig. 14)
    This decomposition facilitates more accurate VLM reasoning.
  • Task completion verification — Once the agent considers the task completed, the VLM utilizes observation images from the past steps to perform a final inference, thereby validating task credibility. (As shown in Fig. 15)
To better understand the process of VLM reasoning, we present the prompts used for the three VLM reasoning scenarios described above.
Figure 13: Prompt 2 for “find target in exploration inference”. The placeholders {…} will be replaced by the corresponding information. The gray-highlighted text represents the information of the full M3DSG.
Figure 14: Prompt 3 for “explore frontier in exploration inference”. The placeholders {…} will be replaced by the corresponding information. The gray-highlighted text represents the information of frontier images as in 3D-Mem [41].
Figure 15: Prompt 4 for “task completion verification”. The placeholders {…} will be replaced by the corresponding information. The gray-highlighted text represents the information of observation images from the past steps.